Please help me solve this equation:
If f(x) = x^2 - 2x + 3, find f(x + 2)

Answer by nerdybill(7384)

f(x) = x^2 - 2x + 3
.
f(x+2) = (x+2)^2 -2(x+2) +3
f(x+2) = (x+2)(x+2) -2(x+2) +3
f(x+2) = x^2+4x+4 -2(x+2) +3
f(x+2) = x^2+4x+4 -2x-2 +3
f(x+2) = x^2+2x+4-2 +3
f(x+2) = x^2+2x+2 +3
f(x+2) = x^2+2x+5